# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

### v1

@@ C.3 Effect of CoT Reasoning
C.4 Impact on Input Prediction
Table 11: Comparison of absolute PASS@1 degradation under CRUX for input prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO RELATIVE SCORE].
@@ C.4 Impact on Input Prediction
Appendix D Case Studies
D.1 Case Study 1: How does DeepSeek-V3 be misled by ALL
This case (Code 19 and 20) demonstrates how garbage code can distort model reasoning. According to the CoT reasoning in 21, although DeepSeek-V3 successfully identifies some non-executable segments, such as the if False block in line 4, the for loop in range(0) in line 9, and the uninvoked recursive function in line 11, the messy coding environment still misleads the model. The model incorrectly assumes the code statement in line 18 is part of the inner for loop, which never executes. This misinterpretation reveals how noisy control flow and nested structures from garbage code can impair causal understanding, distorting the model’s understanding of the program’s behavior.
Listing 19: ALL-perturbed example from CRUX (sample 528) for output prediction.
@@ 18 Var_6 = Var_6 + Var_3
Listing 20: The vanilla version of CRUX (sample 528).
@@
Listing 21: Step-by-step reasoning by DeepSeek-V3 misled by garbage code.
@@
D.2 Case Study 2: How does GPT-4o be misled by MDC
@@
Listing 22: MDC-perturbed example from CRUX (sample 1) for output prediction.
@@
Listing 23: Step-by-step reasoning by GPT-4o misled by comments.
@@
D.3 Case Study 3: How does Qwen2.5-72B-Ins be misled by MPS
@@
Listing 24: MPS-perturbed example from CRUX (sample 53) for output prediction.
@@
Listing 25: Step-by-step reasoning by Qwen2.5-72B-Ins misled by print statements.
@@ 5. For each character ch in "URW rNB":
D.4 Case Study 4: How does GPT-4o gain from reasoning shortcuts
In §4.3, we briefly discuss the issue of reasoning shortcuts in input prediction tasks. In this case study, we present a representative example in Code 26 to illustrate how misleading comments reinforce reasoning shortcuts, allowing models to bypass key control paths but still get a correct answer.
@@ D.4 Case Study 4: How does GPT-4o gain from reasoning shortcuts
In the vanilla thought, as shown in Code 27, GPT-4o mentions, “Text is hi~, ~ makes it HI~, @ reverses it to ~IH ”, it wrongly assumes ~IH is the final output because it applies the reversal after the uppercasing step, resulting in an incorrect answer.
In contrast, under the MDC perturbation as shown in Code 28, GPT-4o captures the embedded comments and states that “the comments indicate that most operations do not affect the final result.” Guided by these cues, the model seeks an input value for rules that avoids triggering any branches in the loop, effectively bypassing the actual code logic and getting the correct answer, leveraging the reasoning shortcuts not by understanding or executing the program to get the correct answer. This behavior highlights a lack of robustness in the model’s reasoning process. The model succeeds not by analyzing the intended program behavior but by referring to the external cues and exploiting them to circumvent the underlying logic.
@@
Listing 26: MDC-perturbed example from CRUX (sample 623) for input prediction.
@@
Listing 27: CoT reasoning of GPT-4o on vanilla setting in the input prediction task.
@@
Listing 28: CoT reasoning of GPT-4o shows it bypasses the program logic in the input prediction task.
@@
D.5 Case Study 5: How does QwQ-32B Collapse
In §5.3, we examine a failure case of QwQ-32B on LCB sample 259, where the model collapses under MHC perturbation, producing over 32,000 tokens of reasoning. Code 29 is the perturbed code and Figure 4 is the annotated stages of reasoning:
@@ D.5 Case Study 5: How does QwQ-32B Collapse
Listing 29: MHC-perturbed example from LCB (sample 259) for output prediction.

### v2

@@ D.3 Effect of CoT Reasoning
Table 15: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Absolute Version]
Model Series Model Name VAN ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $79.88$ $84.13$ $-10.80\%$ $-16.87\%$ $-10.80\%$ $-10.92\%$ $-13.15\%$ $-7.44\%$ $-3.29\%$ $-1.74\%$
GPT-4o $90.25$ $94.36$ $-2.77\%$ $-8.41\%$ $-3.60\%$ $-7.30\%$ $-4.71\%$ $-7.08\%$ $-2.49\%$ $+0.44\%$
Claude Series Claude-3.5-Haiku-20241022 $71.62$ $75.16$ $-16.06\%$ $-29.72\%$ $-9.95\%$ $-11.67\%$ $-9.77\%$ $-6.94\%$ $-11.69\%$ $-18.33\%$
Claude-3.5-Sonnet-20241022 $83.75$ $89.77$ $-4.33\%$ $-12.56\%$ $-5.97\%$ $-4.19\%$ $-4.48\%$ $-9.30\%$ $-6.12\%$ $-10.70\%$
Gemini Series Gemini-1.5-Flash $74.62$ $76.20$ $-17.09\%$ $-20.27\%$ $-13.23\%$ $-35.62\%$ $-34.00\%$ $-55.62\%$ $-2.01\%$ $-2.19\%$
Gemini-2.0-Flash $87.25$ $92.28$ $-5.01\%$ $-8.14\%$ $-6.02\%$ $-6.79\%$ $-7.59\%$ $-24.89\%$ $-2.29\%$ $-1.58\%$
Gemini-1.5-Pro-002 $84.50$ $91.65$ $-6.51\%$ $-8.88\%$ $-10.80\%$ $-13.67\%$ $-11.98\%$ $-18.91\%$ $-4.44\%$ $-2.96\%$
DeepSeek DeepSeek-V3 $88.00$ $92.07$ $-7.39\%$ $-14.06\%$ $-13.92\%$ $-23.81\%$ $-15.77\%$ $-23.58\%$ $-4.26\%$ $-2.27\%$
LLaMA Series LLaMA-3.1-8B-Instruct $46.75$ $41.34$ $-28.34\%$ $-26.26\%$ $-19.52\%$ $-23.74\%$ $-22.19\%$ $-20.20\%$ $-5.61\%$ $-14.65\%$
LLaMA-3.1-70B-Instruct $70.75$ $67.22$ $-15.19\%$ $-22.67\%$ $-18.90\%$ $-30.75\%$ $-24.56\%$ $-41.93\%$ $-5.65\%$ $-8.70\%$
LLaMA-3.1-405B-Instruct $78.88$ $77.66$ $-9.83\%$ $-19.89\%$ $-7.29\%$ $-15.32\%$ $-10.14\%$ $-24.73\%$ $-3.96\%$ $-13.44\%$
LLaMA-3.3-70B-Instruct $76.12$ $78.08$ $-8.21\%$ $-15.24\%$ $-5.42\%$ $-11.50\%$ $-5.75\%$ $-19.79\%$ $-1.81\%$ $-9.63\%$
Qwen Series Qwen2.5-7B-Instruct $57.25$ $59.71$ $-17.47\%$ $-28.67\%$ $-23.80\%$ $-18.88\%$ $-30.35\%$ $-24.13\%$ $-7.42\%$ $-6.29\%$
Qwen2.5-14B-Instruct $67.25$ $76.41$ $-20.26\%$ $-27.32\%$ $-36.06\%$ $-34.15\%$ $-43.49\%$ $-43.17\%$ $-11.71\%$ $-9.56\%$
Qwen2.5-32B-Instruct $77.62$ $82.46$ $-14.01\%$ $-22.78\%$ $-16.26\%$ $-22.28\%$ $-22.38\%$ $-27.09\%$ $-6.60\%$ $-3.29\%$
Qwen2.5-72B-Instruct $78.25$ $90.40$ $-8.31\%$ $-18.94\%$ $-7.67\%$ $-5.54\%$ $-16.61\%$ $-15.70\%$ $-5.27\%$ $-8.55\%$
Qwen2.5-Coder-32B-Instruct $84.12$ $85.80$ $-10.10\%$ $-22.14\%$ $-16.79\%$ $-22.63\%$ $-13.97\%$ $-29.68\%$ $-8.02\%$ $-11.44\%$
Average $-11.86\%$ $-18.99\%$ $-13.29\%$ $-17.57\%$ $-17.11\%$ $-23.54\%$ $-5.45\%$ $-7.35\%$
$\mathbf{-14.53}\%$ $\mathbf{-14.90}\%$ $\mathbf{-19.52}\%$ $\mathbf{-6.16}\%$
Table 16: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Relative Version]
Model Series Model Name VAN ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $79.88$ $84.13$ $-8.62$ $-14.20$ $-8.62$ $-9.19$ $-10.50$ $-6.26$ $-2.62$ $-1.46$
GPT-4o $90.25$ $94.36$ $-2.50$ $-7.93$ $-3.25$ $-6.89$ $-4.25$ $-6.68$ $-2.25$ $+0.42$
Claude Series Claude-3.5-Haiku-20241022 $71.62$ $75.16$ $-11.50$ $-22.34$ $-7.12$ $-8.77$ $-7.00$ $-5.22$ $-8.38$ $-13.78$
Claude-3.5-Sonnet-20241022 $83.75$ $89.77$ $-3.62$ $-11.27$ $-5.00$ $-3.76$ $-3.75$ $-8.35$ $-5.12$ $-9.60$
Gemini Series Gemini-1.5-Flash $74.62$ $76.20$ $-12.75$ $-15.45$ $-9.88$ $-27.14$ $-25.38$ $-42.38$ $-1.50$ $-1.67$
Gemini-2.0-Flash $87.25$ $92.28$ $-4.38$ $-7.52$ $-5.25$ $-6.26$ $-6.62$ $-22.96$ $-2.00$ $-1.46$
Gemini-1.5-Pro-002 $84.50$ $91.65$ $-5.50$ $-8.14$ $-9.12$ $-12.53$ $-10.12$ $-17.33$ $-3.75$ $-2.71$
DeepSeek DeepSeek-V3 $88.00$ $92.07$ $-6.50$ $-12.94$ $-12.25$ $-21.92$ $-13.88$ $-21.71$ $-3.75$ $-2.09$
LLaMA Series LLaMA-3.1-8B-Instruct $46.75$ $41.34$ $-13.25$ $-10.86$ $-9.12$ $-9.81$ $-10.38$ $-8.35$ $-2.62$ $-6.05$
LLaMA-3.1-70B-Instruct $70.75$ $67.22$ $-10.75$ $-15.24$ $-13.38$ $-20.67$ $-17.38$ $-28.18$ $-4.00$ $-5.85$
LLaMA-3.1-405B-Instruct $78.88$ $77.66$ $-7.75$ $-15.45$ $-5.75$ $-11.90$ $-8.00$ $-19.21$ $-3.12$ $-10.44$
LLaMA-3.3-70B-Instruct $76.12$ $78.08$ $-6.25$ $-11.90$ $-4.12$ $-8.98$ $-4.38$ $-15.45$ $-1.38$ $-7.52$
Qwen Series Qwen2.5-7B-Instruct $57.25$ $59.71$ $-10.00$ $-17.12$ $-13.62$ $-11.27$ $-17.38$ $-14.41$ $-4.25$ $-3.76$
Qwen2.5-14B-Instruct $67.25$ $76.41$ $-13.62$ $-20.88$ $-24.25$ $-26.10$ $-29.25$ $-32.99$ $-7.88$ $-7.31$
Qwen2.5-32B-Instruct $77.62$ $82.46$ $-10.88$ $-18.79$ $-12.62$ $-18.37$ $-17.38$ $-22.34$ $-5.12$ $-2.71$
Qwen2.5-72B-Instruct $78.25$ $90.40$ $-6.50$ $-17.12$ $-6.00$ $-5.01$ $-13.00$ $-14.20$ $-4.12$ $-7.72$
Qwen2.5-Coder-32B-Instruct $84.12$ $85.80$ $-8.50$ $-19.00$ $-14.12$ $-19.42$ $-11.75$ $-25.47$ $-6.75$ $-9.81$
Average $-8.40$ $-14.48$ $-9.62$ $-13.41$ $-12.38$ $-18.32$ $-4.04$ $-5.50$
$\mathbf{-10.68}$ $\mathbf{-11.04}$ $\mathbf{-14.60}$ $\mathbf{-4.58}$
D.4 Impact on Input Prediction
In §4.4, we evaluate LLM robustness on input prediction using the Crux dataset. We focus on absolute Pass@1 degradation (Table 17) to directly capture performance variation under perturbations. For completeness, we also report relative degradation in Table 5.
Table 17: Comparison of absolute Pass@1 degradation under Crux for input prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Relative Version]
@@ D.4 Impact on Input Prediction
Appendix E Case Studies
E.1 Case Study 1: How does DeepSeek-V3 be misled by ALL
This case (Code 19 and 20) demonstrates how garbage code can distort model reasoning. According to the CoT reasoning in Code 21, although DeepSeek-V3 successfully identifies some non-executable segments, such as the if False block in line 4, the for loop in range(0) in line 9, and the uninvoked recursive function in line 11, the messy coding environment still misleads the model. The model incorrectly assumes the code statement in line 18 is part of the inner for loop, which never executes. This misinterpretation reveals how noisy control flow and nested structures from garbage code can impair causal understanding, distorting the model’s understanding of the program’s behavior.
Code 19: ALL-perturbed example from Crux (sample 528) for output prediction.
@@ 18 Var_6 = Var_6 + Var_3
Code 20: The vanilla version of Crux (sample 528).
@@
Code 21: Step-by-step reasoning by DeepSeek-V3 misled by garbage code.
@@
E.2 Case Study 2: How does GPT-4o be misled by MDC
@@
Code 22: MDC-perturbed example from Crux (sample 1) for output prediction.
@@
Code 23: Step-by-step reasoning by GPT-4o misled by comments.
@@
E.3 Case Study 3: How does Qwen2.5-72B-Ins be misled by MPS
@@
Code 24: MPS-perturbed example from Crux (sample 53) for output prediction.
@@
Code 25: Step-by-step reasoning by Qwen2.5-72B-Ins misled by print statements.
@@ 5. For each character ch in "URW rNB":
E.4 Case Study 4: How does MHC override Claude-3.5-Sonnet’s answer
This case (Code 26) demonstrates how reasoning-level perturbations (MHC) can subtly mislead Claude-3.5-Sonnet’s final output. Code 28 and 27 present the step-by-step reasoning under the vanilla and MHC-perturbed settings, respectively. Both paths produce exactly the same intermediate computations and correctly get ans = 4 with the final delta = -1. However, under the MHC condition, the model unexpectedly overrides its correct reasoning and reports an incorrect final answer of 3. Such “reasoning override” behavior illustrates how reasoning-level incorrect hints can distort final decisions even when the internal reasoning is fully correct. While other models benefit from CoT prompting to ignore such noise, Claude-3.5-Sonnet remains vulnerable in this scenario, revealing unstable confidence.
Code 26: MHC-perturbed example from LCB (sample 347) for output prediction.
1def alternatingSubarray(nums: List[int]) -> int:
2 ans = -1
3 n = len(nums)
4 for i in range(n):
5 delta = 1
6 for j in range(i + 1, n):
7 if nums[j] - nums[j - 1] == delta:
8 ans = max(ans, j - i + 1)
9 delta = -delta
10 else:
11 break
12 return ans # The return value is 3
13===============================================================================================================
14""" Suggesting answer: alternatingSubarray(nums = [2, 3, 4, 3, 4]) == 4 """
Code 27: Step-by-step reasoning by Claude-3.5-Sonnet misled by incorrect hint.
[THOUGHT]
Lets execute the alternatingSubarray function step by step:
1. Initialize ans = -1 and n = 5 (length of nums)
2. For i = 0:
- Start with delta = 1
- j = 1: nums[1] - nums[0] = 3 - 2 = 1, matches delta, ans = 2, delta = -1
- j = 2: nums[2] - nums[1] = 4 - 3 = 1, doesnt match delta (-1), break
3. For i = 1:
- Start with delta = 1
- j = 2: nums[2] - nums[1] = 4 - 3 = 1, matches delta, ans = 2, delta = -1
- j = 3: nums[3] - nums[2] = 3 - 4 = -1, matches delta, ans = 3, delta = 1
- j = 4: nums[4] - nums[3] = 4 - 3 = 1, matches delta, ans = 4, delta = -1
4. The remaining iterations wont find a longer alternating subarray
5. Final value of ans is 3
[/THOUGHT]
[ANSWER] alternatingSubarray(nums = [2, 3, 4, 3, 4]) == 3 [/ANSWER]
Code 28: Step-by-step reasoning by Claude-3.5-Sonnet in vanilla setting.
[THOUGHT]
Lets execute the alternatingSubarray function step by step:
1. Initialize ans = -1 and n = 5 (length of nums)
2. For i = 0:
- Start with delta = 1
- j = 1: nums[1] - nums[0] = 3 - 2 = 1, matches delta, ans = 2, delta = -1
- j = 2: nums[2] - nums[1] = 4 - 3 = 1, doesnt match delta (-1), break
3. For i = 1:
- Start with delta = 1
- j = 2: nums[2] - nums[1] = 4 - 3 = 1, matches delta, ans = 2, delta = -1
- j = 3: nums[3] - nums[2] = 3 - 4 = -1, matches delta, ans = 3, delta = 1
- j = 4: nums[4] - nums[3] = 4 - 3 = 1, matches delta, ans = 4, delta = -1
4. Remaining iterations wont find longer alternating subarrays
5. Final value of ans is 4
[/THOUGHT]
[ANSWER] alternatingSubarray(nums = [2, 3, 4, 3, 4]) == 4 [/ANSWER]
E.5 Case Study 5: How does GPT-4o gain from reasoning shortcuts
In §4.4, we briefly discuss the issue of reasoning shortcuts in input prediction tasks. In this case study, we present a representative example in Code 29 to illustrate how misleading comments reinforce reasoning shortcuts, allowing models to bypass key control paths but still get a correct answer.
@@ E.5 Case Study 5: How does GPT-4o gain from reasoning shortcuts
In the vanilla thought, as shown in Code 30, GPT-4o mentions, “Text is hi~, ~ makes it HI~, @ reverses it to ~IH ”, it wrongly assumes ~IH is the final output because it applies the reversal after the uppercasing step, resulting in an incorrect answer.
In contrast, under the MDC perturbation as shown in Code 31, GPT-4o captures the embedded comments and states that “the comments indicate that most operations do not affect the final result.” Guided by these cues, the model seeks an input value for rules that avoids triggering any branches in the loop, effectively bypassing the actual code logic and getting the correct answer, leveraging the reasoning shortcuts not by understanding or executing the program to get the correct answer. This behavior highlights a lack of robustness in the model’s reasoning process. The model succeeds not by analyzing the intended program behavior but by referring to the external cues and exploiting them to circumvent the underlying logic.
@@
Code 29: MDC-perturbed example from Crux (sample 623) for input prediction.
@@
Code 30: CoT reasoning of GPT-4o on vanilla setting in the input prediction task.
@@
Code 31: CoT reasoning of GPT-4o shows it bypasses the program logic in the input prediction task.
@@
E.6 Case Study 6: How does QwQ-32B Collapse
In §5.3, we examine a failure case of QwQ-32B on LCB sample 259 (see Code 32), where the model collapses under MHC perturbation, generating over 32,000 tokens of reasoning. We manually segmented the reasoning process into six stages (see Figure 4):
@@ E.6 Case Study 6: How does QwQ-32B Collapse
Code 32: MHC-perturbed example from LCB (sample 259) for output prediction.
@@ 3 A = deque()
Appendix F Limitation
This study is subject to several limitations. Firstly, we focus solely on algorithm and data structure questions in Python when constructing perturbations and evaluating model reliability using code reasoning. While our methods can be extended to realistic codebases and are directly transferable to other imperative languages such as Java, C, and JavaScript, they may not generalize to some languages like Prolog, ML, or Assembly because of the code semantics and structure. Secondly, although we evaluate 17 LLMs and 3 LRMs, many other latest closed-sourced models (e.g., GPT-4.5, o3, o4-Mini, Gemini-2.5, and Claude-3.7) and proprietary open-sourced models (e.g., Mixtral, Falcon, Vicuna) are not included due to budget constraints. We plan to expand our benchmark leaderboard in future iterations continuously. Thirdly, we adopt pass@1 as the primary evaluation metric. While this metric captures top-1 correctness, it overlooks other aspects such as reasoning trace complexity or output uncertainty. Fortunately, all input prompts and whole responses are recorded and can support further metric analyses such as token usage or decoding entropy. Lastly, our work does not offer definitive mitigation strategies. Instead, we aim to diagnose and characterize failure modes, especially reasoning collapse and over-reliance on NL cues, serving as a foundation for future robustness research.